OpenStudy (anonymous):

How do you solve for the discriminant of a quadrant? Ex: 3m^2+8m-4=3 Choices: A)148 B)-47 C)112 D)-20

OpenStudy (anonymous):

\[Discriminant = b^2-4 \times a \times c\]

OpenStudy (anonymous):

\[3m^2+8m-4=3\] \[3m^2+8m-7=0\] a=3 , b=8 , c=-7

OpenStudy (anonymous):

Now that you know what a, b, and c are equal to, plug it in the formula above.

OpenStudy (anonymous):

Where did the 4 in the first equation come from?

OpenStudy (anonymous):

4 is part of the quadratic formula, it's always there.

OpenStudy (anonymous):

Okay Thank you